A scale shows the weight of a baby elephant t0 be 150.,pounds,,J,t,ie'scale ,.. has an accuracy of plus or minus 3 pounds. Let the true weight oMne · ' elephant be w. Write an absolute value inequality to show that w is within 3 pounds of 150 pounds. Then solve the inequality. What is the range for the true weight of the baby elephant?
Answer by ikleyn(52799)

.

The absolute value inequality, which you want to get, is THIS ONE

    |x - 150| <= 3  pounds.


Its solution, step by step, is as follows

    -3 <= x - 150 <= 3

    150-3 <= x <= 150+3

      147 <= x <=   153.

Solved, with explanations.